# Trade-offs between structural richness and communication efficiency in music network representations

### v2

@@ Abstract
Music is a structured and perceptually rich sequence of sounds in time with well-defined symbolic features, whose perception is shaped by the interplay of expectation and uncertainty. Network science offers a powerful framework for studying its structural organization and communication efficiency. However, it remains unclear how feature selection affects the properties of reconstructed networks and perceptual alignment. Here, we systematically compare eight encodings of musical sequences, ranging from single-feature descriptions to richer multi-feature combinations. We show that representational choices fundamentally shape network topology, the distribution of uncertainty, and the estimated communication efficiency under perceptual constraints. Single-feature representations compress sequences into dense transition structures that support efficient communication, yielding high entropy rates with low modeled perceptual error, but they discard structural richness. By contrast, multi-feature representations preserve descriptive detail and structural specificity, expanding the state space and producing sharper transition profiles and lower entropy rates, which leads to higher modeled perceptual error. Across representations, we found that uncertainty increasingly concentrates in nodes with higher diffusion-based centrality while their perceptual error remains low, unveiling an interplay between predictable structure and localized surprise. Together, these results show that feature choice directly shapes music network representation, describing trade-offs between descriptive richness and communication efficiency and suggesting structural conditions that may support efficient learning and prediction.
@@ I Introduction
Music is a highly complex phenomenon built from low-level features that combine into patterns incorporating multiple structural aspects. Gaining insight into how these patterns emerge from basic musical events offers a useful perspective for understanding both the complexity of music and the perceptual processes involved in listening Koelsch (2014); Koelsch et al. (2019), while also providing a model system for studying how complex structure arises in sequential data more generally.
@@
Musical sequences can be represented at different levels of detail by selecting subsets of features—such as pitch, interval, or duration—that preserve some distinctions while discarding others, an idea closely related to viewpoint-based approaches in music cognition Conklin and Witten (1995). To study how these representational choices shape the resulting musical structure, network representations can be employed, in which sequence elements are modeled as nodes and observed transitions between elements as edges. Prior work used pitch-based representations with octave and duration Liu et al. (2009); Ferretti (2017), pitch-class sets Gomez et al. (2014); serrà_corral_boguñá_haro_arcos_2012; Nardelli (2023); Manaris et al. (2005); Serra-Peralta et al. (2021), and chord-split representations Kulkarni et al. (2024); Di Marco et al. (2025), which resulted in different topologies and statistical insights del Río et al. (2008); Zanette (2006); Moss et al. (2019); Mehr et al. (2019); Rohrmeier and Cross (2008). Understanding how these representational choices alter network topology and information structure requires a systematic comparison across levels of detail, one that makes explicit the trade-off between complexity reduction and the retention of relevant structure Tishby et al. (2000).
@@
A second key aspect of representational choice concerns not the amount of information detail it preserves, but how its organization alters the uncertainty in musical sequence continuation that humans perceive. In particular, recent work shows that communication efficiency under perceptual constraints–quantified by comparing the true uncertainty in a sequence with the uncertainty humans effectively exhibit–depends on how that uncertainty is organized Lynn et al. (2020a). Human responses to sequential input are limited by memory, noise, and representational cost Lynn et al. (2020b); Momennejad et al. (2017); Howard and Kahana (2002); Meyniel et al. (2016), so the way transitions are organized directly affects how easily the sequence can be processed. For instance, Lynn et al. Lynn et al. (2020b) showed, using reaction-time experiments with visual sequences, that even when two sequences have the same average uncertainty, differences in the organization of transitions can cause humans to react as if the uncertainty were higher, resulting in a perceptual error—a mismatch between the true transition statistics and the observer’s inferred model. That is, structural organization affects the amount of information effectively perceived.
Also, real networks across domains, such as word networks or citation networks, show consistently lower modeled perceptual error than their randomized counterparts, indicating that their transitions are organized in ways that improve communication efficiency Lynn et al. (2020b, a). Crucially, this efficiency seems to differ across representations within the same domain: networks built from content words exhibit lower modeled perceptual error than those built from grammatical words Lynn et al. (2020b), reflecting that representational choices shape transition regularities. Together, these findings suggest that each representation defines its own uncertainty landscape and the way a representation structures uncertainty has an influence on the information available to a listener, and thus implications for perceptual difficulty—precisely the question we address in this work.
@@
Here, we systematically compare eight network representations of music that are drawn from and extend existing approaches in the literature. These vary in their level of detail, from single-feature encodings (pitch, interval, duration) to multi-feature combinations that incorporate octave or duration with pitch. We quantify how these representational choices reshape network topology and evaluate how efficiently each representation structures information by computing both the resulting entropy rate and the divergence from a perceptual-constraint estimate. Together, these measures locate each encoding along a trade-off between descriptiveness—the degree to which the underlying musical structure is preserved—and communication efficiency, the amount of information conveyed with minimal modeled perceptual error Tishby et al. (2000); Lynn et al. (2020b). In addition, by further examining how uncertainty and modeled error are distributed locally across each network representation—and how these patterns vary with piece length—we identify conditions under which random-walk centrality aligns with uncertainty. Such structures may facilitate downstream statistical compression or predictive abstraction Huron (2008); Pearce (2018), offering clues about why certain structural forms are easier for humans to learn. More broadly, this network-based framework links the information-theoretic costs of a representational choice to the structural patterns it reveals, offering a general approach for analyzing how complexity is organized across different representational levels and suggesting structural conditions that may support efficient learning and prediction.
@@ II Results
To examine how different levels of musical description shape structural and informational properties, we constructed eight network representations of the same sequences, spanning both established encodings from the literature and controlled feature-level variants introduced for systematic comparison. We start with minimal, single-feature encodings that expose the basic building blocks of musical structure, and then introduce additional features stepwise, allowing us to observe how topological and informational patterns emerge or transform with increasing representational richness. In all cases, nodes correspond to unique feature configurations and edges record empirical transitions between consecutive elements in the musical sequence. The eight representations are defined as follows:
@@ II.3 Alignment of uncertainty and inference across musical representations
Figure 4: Global alignment of uncertainty and inference in music networks. (a) Median entropy $S$ versus mean node-level entropy $\bar{S}$. (b) Median KL divergence $D_{KL}$ versus node-averaged KL divergence $\bar{D}_{KL}$. (c) Mean node-level entropy with and without edge weights ($\bar{S}$ versus $\bar{S}^{\mathrm{uw}}$). (d) Median node-averaged KL divergence with and without edge weights ($\bar{D}_{KL}$ versus $\bar{D}^{\mathrm{uw}}_{KL}$). (d) Median entropy $S$ versus the difference between median and node-level entropy ($S-\bar{S}$). (e) Median KL divergence $D_{KL}$ vs the difference between median and node-averaged KL divergence ($D_{KL}-\bar{D}_{KL}$). In (d) and (e) data is grouped by transition count. Transition counts are divided into five equal-sized bins. Smaller and more transparent markers correspond to shorter pieces. Points show medians, and error bars indicate the interquartile range. All panels correspond to right-hand tracks. Models are grouped by the number of musical features used. Each symbol corresponds to an individual model.
@@ III Discussion
Representational choice plays a decisive role in the structural properties of musical networks. Our results show that the number and grouping of features—and whether musical events are split or not—substantially influence the resulting topology. Properties often associated with musical corpora, such as small-world organization Ferretti (2017); Liu et al. (2009), heavy-tailed degree distributions, or Zipf- and Heaps-like scaling Manaris et al. (2005); Zanette (2006), emerge only under certain representational levels and are not inherent to the raw data alone. More detailed, multi-feature representations tend to produce markedly heterogeneous and skewed structures, whereas compressed vocabularies attenuate these patterns. These findings show that topological regularities in music depend on how the underlying symbolic space is defined.
Building on this, the representational differences instantiate a broader trade-off between communication efficiency and descriptive richness. The limited perceptual error associated with compressed representations despite their high entropy implicates that information from higher-order transitions several steps apart in the musical sequence do not disturb substantially efficient communication, whereas higher-order transitions in richer representations capturing finer musical structure deviate more importantly from local transitions, which increases perceptual error and lowers entropy rate. This echoes principles from information bottleneck theory Tishby et al. (2000), minimum-description-length formulations Rissanen (1983), and probabilistic models of expectation Pearce and Wiggins (2012), where representations balance fidelity with cognitive cost. Within the network formalism, these trade-offs become explicit: the vocabulary used to encode a sequence not only determines the information carried per step but also its structural cost, captured by the KL divergence between the true transition matrix and the cognitively constrained approximation. Higher KL values indicate that expanded vocabularies introduce statistical structure that is harder for a human-like perceptual system to track, reflecting the additional cost imposed by finer-grained distinctions. By making the structure of information explicit—locally through uncertainty and globally through the stationary distribution and KL divergence–this framework offers a new perspective on how representational choices and perceptual limitations jointly constrain sequential processing.
@@
Although our analysis remains descriptive rather than cognitive, the emergence of efficient communication channels in simpler representations parallels several known properties of auditory and musical processing. Neural pathways handle pitch, duration, and timbre through partially distinct streams Peretz and Coltheart (2003), and hierarchical integration Peretz and Zatorre (2005); Koelsch (2011) resembles the progressive increase in structural complexity we observe as more features are added. At the same time, human expectations depend on learning and exposure Pearce (2018); Daikoku et al. (2021), and thus cannot be reduced to the statistical transitions in a corpus. Our results, therefore, do not model cognition directly, but they provide a structural account of how musical information is organized at different levels of descriptiveness—an organization that perceptual systems must interact with, regardless of the specific statistical learning models involved.
@@
This structural organization also offers insight into why some musical systems support pattern discovery more effectively than others. In richer representations, uncertainty concentrates in dynamically central nodes while less central nodes remain relatively predictable. Such landscapes provide learners with stable contexts and informative deviations, a contrast to uniformly uncertain or near-random systems in which motifs and regularities become difficult to extract. Notably, this interplay between predictable structure and localized surprise has also been linked to affective responses in music, where heterogeneous uncertainty can enhance engagement and pleasure Huron (2008); Cheung et al. (2019); Gold et al. (2019). Although learning is beyond the scope of this work, the results suggest that differentiated uncertainty may facilitate both pattern extraction and the perception of structure in musical sequences.
@@
Our representational hierarchy was chosen to build on established symbolic viewpoints in music-network research while enabling a controlled progression from coarse to detailed descriptions. Other musically informed viewpoints—such as scale-relative functions or learned feature spaces—may yield different structural signatures and should be explored in future work. Similarly, our focus on classical piano music provided a rich context for examining how complexity grows with representational detail, but the same framework can be applied to other repertoires. Our intuition is that genres with more repetitive or minimalist structures will show weaker differentiation across representational levels, offering a useful contrast to the dynamics observed here.
A natural extension of this work is to examine how expectations derived from statistical-learning models interact with the informational landscapes uncovered here. Models such as IDyOM Pearce and Wiggins (2012) introduce learned priors that may bias the inferred transition structure away from the raw distribution of events, effectively modifying perceived topology and uncertainty. Understanding how information at different representational levels contributes to these expectations—and how listeners integrate or prioritize these layers—could provide deeper insight into the efficient processing of musical complexity.
@@
More broadly, this approach demonstrates how the structural and informational consequences of representational choice can be analyzed in any sequential domain. Music serves as a clear testbed because its feature combinations and hierarchical abstractions are well characterized, but the framework is easily transferable to other sequences and systems. By tracing how topologies and informational landscapes evolve with representational detail, we gain a principled way to compare how different feature sets shape the organization of sequential information and the emergence of efficient heterogeneous structures.

### v3

@@ Abstract
Music is a structured and perceptually rich sequence of sounds in time, whose perception is shaped by the interplay of expectation and uncertainty about what comes next. Yet the uncertainty we infer from music depends on how the musical piece is encoded as an event sequence. In this work, we use network representations, in which event types are nodes and observed transitions are directed edges, to compare how different feature encodings shape the transition structure we recover and what this implies for both the descriptive uncertainty expectation under imperfect memory and noise. We systematically analyse eight encodings of piano music, from single-feature vocabularies to richer multi-feature combinations. These representational choices reorganize the state space and fundamentally reshape network topology, shifting how uncertainty is distributed across transitions. To connect these descriptive differences to perception, we adopt a perceptual-constraint model that captures imperfect access to transition statistics. Overall, compressed single-feature representations yield dense transition structures with higher entropy rates, corresponding to higher average uncertainty per step, yet low model error, indicating that the constrained estimate stays close to the corpus transitions. In contrast, richer multi-feature representations preserve finer distinctions but expand the state space, sharpen transition profiles, lower entropy rates, and increase model error. Finally, across representations, uncertainty concentrates in diffusion-central nodes while model error remains low there, suggesting an informational landscape in which predictable flow coexists with localized surprise. Overall, our results show that feature choice shapes not only the networks we reconstruct, but also whether their resulting uncertainty is a plausible proxy for the expectations listeners can realistically learn and use.
@@ I Introduction
Music is a highly complex phenomenon composed of low-level features that combine into patterns that incorporate multiple structural aspects. Gaining insight into how these patterns emerge from basic musical events offers a useful perspective for understanding both the complexity of music and the perceptual processes involved in listening Koelsch (2014); Koelsch et al. (2019). These perceptual processes include the formation of expectations about what comes next and the experience of varying uncertainty and surprise as those expectations are confirmed or violated Koelsch et al. (2019); Cheung et al. (2019); Mas-Herrero and Marco-Pallarés (2025).
Any analysis of musical structure depends on how the musical piece is encoded as a sequence of events. Musical sequences can be represented at different levels of detail by selecting subsets of features—such as pitch, interval, or duration—that preserve some distinctions while discarding others, an idea closely related to viewpoint-based approaches in music cognition Conklin and Witten (1995). These representational choices define the event vocabulary and its granularity, and therefore shape which regularities can be expressed or detected, including uncertainty about what comes next. Accordingly, what matters is not only the overall level of this uncertainty, but also how it is organized in a piece Koelsch et al. (2019); Cheung et al. (2019). Since both depend on the chosen event vocabulary, representational choices can systematically reshape the uncertainty patterns we infer.
For example, consider a short melody of eight notes: (A4–C5–A5–F5–A6–C4–F4–A7). If we encode each note as a distinct event by combining pitch class and octave, the sequence contains eight different event types, so each event type tends to have only one observed next-event option, making transitions appear highly predictable. If, instead, we collapse octave and keep only pitch class, the same melody reduces to the pitch-class sequence (A–C–A–F–A–C–F–A) and to just three event types (A, C, and F) repeated across the eight notes. Each event type can then be followed by different events (e.g., A can be followed by either C or F), increasing uncertainty. Thus, although both encodings describe the same musical surface, they induce different transition statistics, raising a basic methodological question: how does representational detail shape the structure we infer from music?
@@
One way to make such differences explicit is to use network representations, in which sequence elements are modeled as nodes and observed transitions between elements as directed edges. Previous work used pitch-based representations with octave and duration Liu et al. (2009); Ferretti (2017), pitch-class sets Gomez et al. (2014); serrà_corral_boguñá_haro_arcos_2012; Nardelli (2023); Manaris et al. (2005); Serra-Peralta et al. (2021), and chord-split representations Kulkarni et al. (2024); Di Marco et al. (2025), which resulted in different topologies and statistical insights del Río et al. (2008); Zanette (2006); Moss et al. (2019); Mehr et al. (2019); Rohrmeier and Cross (2008). Yet, without a systematic comparison across levels of detail within the same musical corpus, it remains unclear whether the reported topological patterns reflect the music itself or the encoding, motivating a cross-representation analysis that makes explicit how different event vocabularies redistribute uncertainty across the network.
Network representations allow us to describe how different encodings redistribute uncertainty. However, because expectation and surprise are perceptual phenomena, whether these descriptive uncertainties can serve as proxies for listeners’ expectations depends on how reliably the corresponding transition structure can be learned under imperfect memory and noise.
@@
This raises a second question: under what conditions does a representation’s transition structure yield expectations that humans can realistically learn and use under perceptual constraints? Human responses to sequential input are limited by memory, noise, and representational cost Lynn et al. (2020b, a); Momennejad et al. (2017); Howard and Kahana (2002); Meyniel et al. (2016). A key consequence is that observers may not rely only on the immediately observed one-step transitions, but may partially integrate information over multiple steps—treating events that are reachable in two or more steps as if they were more directly available than the true statistics imply. Such integration can cause listeners to treat events that occur only after several steps as if they were possible next events, thereby changing the transition probabilities they infer. Recent work shows that communication efficiency under such perceptual constraints depends on the organization of uncertainty Lynn et al. (2020b). Lynn et al. Lynn et al. (2020a) showed, using reaction-time experiments with visual sequences, that even when two sequences have the same average uncertainty, meaning that continuations are equally unpredictable, differences in the organization of transitions can cause humans to react as if the uncertainty were higher, resulting in a perceptual error, defined as the mismatch between the true transition statistics and the observer’s inferred model.
[…]
In short, representation choice fixes the transition statistics and therefore the distribution of uncertainty. Representing these statistics as a network makes their topology explicit. Because humans do not access transition statistics perfectly, perceptual constraints can make the same statistical uncertainty appear higher or lower, depending on how transitions are organized. This motivates (i) a systematic comparison of network structure across levels of representation detail and (ii) an analysis of how the resulting network topology relates to perceptual efficiency.
@@
In this work, we address this issue by analysing eight network representations of music, drawing on and extending existing approaches in the literature. These vary in their level of detail, from single-feature encodings (pitch, interval, duration) to multi-feature combinations that incorporate octave or duration with pitch. We quantify how these representational choices reshape network topology and evaluate how efficiently each representation structures information by computing both the resulting entropy rate and the divergence from a perceptual-constraint estimate. Together, these measures make explicit a trade-off: increasing representational detail can preserve finer aspects of musical structure, but may produce a transition structure that is harder for humans to learn and use, while simpler encodings may sacrifice detail yet yield more perceptually efficient uncertainty estimates Tishby et al. (2000); Lynn et al. (2020b). In addition, by further examining how uncertainty and modeled error are distributed locally across each network representation—and how these patterns vary with piece length—we identify conditions under which random-walk centrality aligns with uncertainty. By quantifying how topology drives divergences between true transition statistics and a perceptually constrained estimate, our approach both reveals a representation-dependent bias that ideal-access expectation models can overlook Pearce and Wiggins (2012); Pearce (2018) and points toward incorporating realistic memory limits when relating uncertainty to neural Mas-Herrero and Marco-Pallarés (2025) and behavioral data Gold et al. (2019).
@@ II Results
To examine how different levels of musical description shape structural and informational properties, we constructed eight network representations of the same sequences, spanning both established encodings from the literature and controlled feature-level variants introduced for systematic comparison. We start with minimal, single-feature encodings that expose the basic building blocks of musical structure, and then introduce additional features stepwise, allowing us to observe how topological and informational patterns emerge or transform as representational richness increases. In all cases, nodes correspond to unique feature configurations and edges record empirical transitions between consecutive elements in the musical sequence. The eight representations are defined as follows:
@@ II.3 Alignment of uncertainty and inference across musical representations
Figure 4: Global alignment of uncertainty and inference in music networks. (a) Median entropy $S$ versus mean node-level entropy $\bar{S}$. (b) Median KL divergence $D_{KL}$ versus node-averaged KL divergence $\bar{D}_{KL}$. (c) Mean node-level entropy with and without edge weights ($\bar{S}$ versus $\bar{S}^{\mathrm{uw}}$). (d) Median node-averaged KL divergence with and without edge weights ($\bar{D}_{KL}$ versus $\bar{D}^{\mathrm{uw}}_{KL}$). (d) Median entropy $S$ versus the difference between median and node-level entropy ($S-\bar{S}$). (e) Median KL divergence $D_{KL}$ vs the difference between median and node-averaged KL divergence ($D_{KL}-\bar{D}_{KL}$). In (e) and (f) data is grouped by transition count. Transition counts are divided into five equal-sized bins. Smaller and more transparent markers correspond to shorter pieces. Points show medians, and error bars indicate the interquartile range. All panels correspond to right-hand tracks. Models are grouped by the number of musical features used. Each symbol corresponds to an individual model.
@@ III Discussion
Representational choice plays a decisive role in the structural properties of musical networks. Our results show that the number and grouping of features—and whether musical events are split or not—substantially influence the resulting topology. Properties often associated with musical corpora, such as small-world organization Ferretti (2017); Liu et al. (2009), heavy-tailed degree distributions, or Zipf- and Heaps-like scaling Manaris et al. (2005); Zanette (2006), emerge only under certain representational levels and are not inherent to the raw data alone. More detailed, multi-feature representations tend to produce markedly heterogeneous and skewed structures, whereas compressed vocabularies attenuate these patterns. These findings show that topological regularities in music depend on the definition of the underlying symbolic space.
Building on this, our results frame representational differences as a trade-off between descriptive richness and communication efficiency under perceptual constraints. Here, descriptive richness denotes the number of musically relevant distinctions preserved by an encoding. This determines how large and how fine-grained its event vocabulary (state space) is. A richer encoding typically makes one-step transitions more specific and more locally predictable, i.e., it lowers uncertainty. However, it can also make the transition structure harder to learn reliably. By contrast, communication efficiency refers to how well a listener can recover and use the resulting transition statistics. This assumes the listener has limited memory and only noisy access to the sequence. Compressed representations tend to yield lower perceptual error, suggesting that their transition topology can be tracked relatively well even with limited access to transition statistics. By contrast, richer representations that preserve finer details reshape the state space and introduce more complex transition structures, increasing the mismatch between true and constrained transition estimates and, in turn, reducing communication efficiency. This echoes principles from information bottleneck theory Tishby et al. (2000) and minimum-description-length formulations Rissanen (1983), in which representations balance fidelity against cognitive cost. In our network framework, this cost becomes explicit: the representation choice determines not only the information carried per step (entropy rate) but also the structural accessibility of that information, quantified by the divergence between the true transition matrix and its cognitively constrained approximation. Viewed this way, representational choices and perceptual limitations jointly determine the uncertainty a listener effectively experiences.
@@
Although our analysis remains descriptive rather than cognitive, the emergence of efficient communication channels in simpler representations parallels several known properties of auditory and musical processing. Neural pathways handle pitch, duration, and timbre through partially distinct streams Peretz and Coltheart (2003), and hierarchical integration Peretz and Zatorre (2005); Koelsch (2011) resembles the progressive increase in structural complexity we observe as more features are added. At the same time, human expectations depend on learning and exposure Pearce (2018); Daikoku et al. (2021), and thus cannot be reduced to the statistical transitions of the piece under study. Our results, therefore, do not model cognition directly, but they provide a structural account of how musical information is organized and of the cost of accessing it. State-of-the-art models of these musical expectations Pearce and Wiggins (2012) estimate uncertainty directly from learned transitions, implicitly assuming ideal access to prior musical exposure in memory. In this work, we use network representations to make explicit how imperfect access to transition statistics reshapes uncertainty, and to quantify how this effect depends on representational choice by measuring the topology-dependent gap between true and constrained transition estimates. Therefore, a natural extension of this work is to examine how expectations derived from these statistical-learning models interact with the informational landscapes uncovered here.
@@
This structural organization also offers insight into why some musical systems support pattern discovery more effectively than others Di Marco et al. (2025). In richer representations, uncertainty concentrates in dynamically central nodes, whereas less central nodes remain relatively predictable. Such landscapes provide learners with stable contexts and informative deviations, a contrast to uniformly uncertain or near-random systems in which motifs and regularities become difficult to extract. Notably, this interplay between predictable structure and localized surprise has also been linked to affective responses in music, where heterogeneous uncertainty can enhance engagement and pleasure Huron (2008); Cheung et al. (2019); Gold et al. (2019). Although learning is beyond the scope of this work, the results suggest that differentiated uncertainty may facilitate both pattern extraction and the perception of structure in musical sequences.
@@
Our representational hierarchy was chosen to build on established symbolic viewpoints in music-network research while enabling a controlled progression from coarse to detailed descriptions. Other musically informed viewpoints—such as tonality-relative encodings based on scale degree or harmonic function (e.g., tonic/dominant), or learned feature spaces—may yield different structural signatures and should be explored in future work. Similarly, our focus on classical piano music provided a rich context for examining how complexity grows with representational detail, but the same framework can be applied to other repertoires. Our intuition is that genres with more repetitive or minimalist structures will show weaker differentiation across representational levels, offering a useful contrast to the dynamics observed here.
@@
More broadly, this approach demonstrates how the structural and informational consequences of representational choice can be analyzed in any sequential domain. Music serves as a clear testbed because its feature combinations and hierarchical abstractions are well-characterized, yet the framework is readily transferable to other sequences and systems. By tracing how topologies and informational landscapes evolve with representational detail, we gain a principled way to compare how different feature sets shape the organization of sequential information and the emergence of efficient heterogeneous structures.